# Epidemiology-informed Graph Neural Network for Heterogeneity-aware Epidemic Forecasting

### v1

@@ III-B Epidemiology-informed Embedding Learning Module
The STGL module only captures the spatio-temporal heterogeneity among different locations, while many previous approaches predict epidemic transmission directly based on the learned ST embedding. However, only capturing the spatio-temporal pattern of the infected population is noisy for epidemic forecasting and lacks the intrinsic epidemic transmission mechanism, which is prone to producing generalization errors in predictions[18]. Therefore, we use the epidemiology mechanistic model to enhance the process of learning location transmission dynamics to capture latent patterns and reflect well-defined disease dissemination mechanisms. Given a location $i$, we assign the time-varying SIR model as the epidemiology mechanistic model, which can be described as following ordinary differential equations (ODEs):
@@
where $S_{i}$, $I_{i}$, and $R_{i}$ represent the number of susceptible, infected, and recovered populations in location $i$, respectively. $\beta_{i}$ and $\gamma_{i}$ are the time-dependent parameters, which mean the infection rate of susceptible populations and the recovery rate of infected populations in the location $i$. Here, we follow the assumption that the total population $N_{i}=S_{i}+I_{i}+R_{i}$ consists of susceptible, infected, and recovered populations.
Although the time-varying SIR model reflects the intrinsic epidemic transmission mechanism, it requires high computational costs and domain knowledge to predict its large-scale parameters simultaneously. In addition, the variables of the time-varying SIR model are heavily entangled, and it is difficult to focus only on predicting changes in the number of infected populations ($I$) in epidemic forecasting. These challenges motivate us to propose the EIEL module that uses node embeddings to parameterize the five time-varying SIR variables. The EIEL module uses five different multi-layer perceptrons (MLPs) to predict susceptible $S$, infected $I$, recovered $R$, infection rate $\beta$, and recovery rate $\gamma$ in the time-varying SIR model at each location and time step. Each MLP consists of a fully connected (FC) layer with $L$ layers and an output layer. Additionally, to enhance the embedding capacity of the EIEL module, we learn epidemiology-informed location embeddings based on the ST embedding $\boldsymbol{P}_{t}=\{\boldsymbol{p}_{t,1},\boldsymbol{p}_{t,2},\ldots,\boldsymbol{p}_{t,N}\}\in\mathbb{R}^{N\times D_{1}}$. Then, based on $i$-th location ST embedding $\boldsymbol{p}_{t,i}\in\mathbb{R}^{D_{1}}$, the epidemiology-informed location embedding $\boldsymbol{e}_{j,i}\in\mathbb{R}^{D_{2}}$ of $j$-th time-varying SIR variable in $i$-th location can be computed by the $j$-th FC layer with $L$ layers:
$\boldsymbol{e}_{j,i}=g_{j}\left(\boldsymbol{p}_{t,i};\{\boldsymbol{W}^{(l)}_{j},\boldsymbol{b}^{(l)}_{j},\sigma^{(l)}_{j}\}_{l=1}^{L}\right),$ (4)
where $\boldsymbol{W}_{j}^{(l)}$ is the weight matrix of the $l$-th hidden layer of the $j$-th FC layer, $\boldsymbol{b}_{j}^{(l)}$ is its bias vector, and $\sigma_{j}^{(l)}$ is the activation function. Through five FC layers, we obtain the epidemiology-informed location embeddings $\boldsymbol{h}_{t+h,i}=[\boldsymbol{e}_{1,i};\boldsymbol{e}_{2,i};\boldsymbol{e}_{3,i};\boldsymbol{e}_{4,i};\boldsymbol{e}_{5,i}]\in\mathcal{R}^{5D_{2}}$ for $i$-th location in time step $t+h$. To predict the dynamic epidemic transmission, the output of MLP is defined as follows:
$y_{j,i}=\sigma_{j}^{(2)}(\boldsymbol{W}_{j}^{(2)}\boldsymbol{e}_{j,i}+\boldsymbol{b}_{j}^{(2)}),$ (5)
where $y_{j,i}\in\{\hat{S}_{t+h,i},\hat{I}_{t+h,i},\hat{R}_{t+h,i},\hat{\beta}_{t+h,i},\hat{\gamma}_{t+h,i}\}$ one of five variables of the susceptible $S$, infected $I$, recovered $R$, infection rate $\beta$, and recovery rate $\gamma$ at location $i$ at time step $t+h$. $\boldsymbol{W}_{j}^{(2)}$, $\boldsymbol{b}_{j}^{(2)}$, and $\sigma_{j}^{(2)}$ are the weight matrix, bias vector, and activation function in the output layer of the $j$-th MLP, respectively. In Section III-E, we introduce the optimization of the EIEL module to ensure that the learned epidemiology-informed location embedding is consistent with the intrinsic epidemic transmission mechanism.
III-C Transmission Graph Module
III-C1 Mechanistic Affinity Graph
The EIEL captures the intrinsic epidemic transmission mechanism of each location separately but cannot learn the interactive relationships between them. To further capture the differences in intrinsic epidemic transmission mechanisms between locations, we propose a time-varying mechanistic affinity graph (MAG) that can efficiently characterize the mechanistic heterogeneity between locations. The MAG $\widetilde{\boldsymbol{M}}_{t}\in\mathbb{R}^{N\times N}$ is calculated by the cosine similarity function $\phi$ between the epidemiology-informed location embeddings of different locations. The mechanistic similarity between locations is defined as follows:
$m_{ij}=\phi(\boldsymbol{h}_{t,i},\boldsymbol{h}_{t,j}),$ (6)
where the mechanistic similarity $m_{ij}\in\boldsymbol{M}_{t}$ represent the mechanistic dependence between $i$-th location and $j$-th location based on epidemiology-informed location embeddings ($\boldsymbol{h}_{t,i}$ and $\boldsymbol{h}_{t,j}$) of two locations at time step $t$. $\boldsymbol{M}_{t}$ is a symmetric matrix. Then, we normalize the similarity between different locations:
$\widetilde{\boldsymbol{M}}_{t}=\boldsymbol{D}^{-\frac{1}{2}}\boldsymbol{M}_{t}\boldsymbol{D}^{-\frac{1}{2}},$ (7)
where $\boldsymbol{D}$ is degree matrix of $\boldsymbol{M}_{t}$. To focus on the key mechanistic dependence, reduce redundant information, and enhance the computational efficiency of GNNs, we design an indicator function to sparsify the MAG $\widetilde{\boldsymbol{M}}_{t}\in[0,1]$. This sparsification is controlled by a sparse threshold $\delta$, applied to each normalized similarity $\widetilde{m}_{ij}\in\widetilde{\boldsymbol{M}}_{t}$:
$\widetilde{m}_{ij}=\begin{cases}\widetilde{m}_{ij}&\text{ if }\widetilde{m}_{ij}\geq\delta\\ 0&\text{ otherwise. }\end{cases},$ (8)
III-C2 Heterogeneous Transmission Graph Network
@@
To fully capture the complex mechanistic heterogeneity by utilizing the mechanistic dependence in the MAG, we propose the Heterogeneous Transmission Graph Network (HTGN) to encode the transmission embedding between locations. The HTGN integrates a MAG with the architecture of Spatio-Temporal Graph Convolutional Networks (STGCN) [27], which comprises two spatio-temporal convolutional blocks (ST-blocks). Given that the historical epidemic transmission risk $\boldsymbol{X}_{t-w+1:t}$ and a MAG $\widetilde{\boldsymbol{M}}_{t}$ in time step $t$, we learn the transmission embedding $\boldsymbol{T}_{t}\in\mathbb{R}^{N\times D_{3}}$ of mechanistic heterogeneity by ST-Blocks:
$\boldsymbol{T}_{t}=\text{ST-Blocks}(\widetilde{\boldsymbol{M}}_{t},\boldsymbol{X}_{t-w+1:t}),$ (9)
III-D Epidemic Forecasting Decoder Module
@@
Finally, by simultaneously capturing the spatio-temporal heterogeneity and mechanistic heterogeneity across different locations, our framework applies an FC layer to predict future epidemic transmission risk, providing the epidemic forecasting result $\hat{\boldsymbol{Y}}_{t+h}$ in time step $t+h$:
$\hat{\boldsymbol{Y}}_{t+h}=\boldsymbol{W}[\boldsymbol{P}_{t};\boldsymbol{T}_{t}]+\boldsymbol{b},$ (10)
@@
where $\boldsymbol{W}$ and $\boldsymbol{b}$ are the weight matrix and basis vector, respectively, $\boldsymbol{P}_{t}$ is the ST embedding learned from the STGL module and $\boldsymbol{T}_{t}$ is the transmission embedding learned from the G module.
III-E Optimization
@@
To ensure the epidemiology-informed location embeddings are consistent with the mechanism of the intrinsic epidemic transmission, we use the idea from the PINN [28] to propose a physics loss to inform the learning process in the EIEL module. The physics loss incorporates the underlying physical laws of the intrinsic epidemic transmission mechanism into the EIEL module. Specifically, we aim to solve the time-varying SIR model using five MLPs by minimizing the multi-objective physics loss $\mathcal{L}_{p}$. Given the prediction of five variables ($\hat{S}_{t+h,i}$, $\hat{I}_{t+h,i}$, $\hat{R}_{t+h,i}$, $\hat{\beta}_{t+h,i}$, and $\hat{\gamma}_{t+h,i}$) of time-varying SIR model, the physics loss $\mathcal{L}_{p}$ to be described as follows:
$\begin{gathered}\mathcal{L}_{d}(\hat{I}_{t+h,i})=\frac{1}{N}\sum_{i=1}^{N}|\hat{I}_{t+h,i}-x_{t+h,i}|,\\ \mathcal{L}_{o}(\hat{S}_{t+h,i},\hat{I}_{t+h,i},\hat{\beta}_{t+h,i},\hat{\gamma}_{t+h,i})\\ =\frac{1}{N}\sum_{i=1}^{N}|\frac{d\hat{I}_{t+h,i}}{d}-(\hat{\beta}_{t+h,i}\hat{S}_{t+h,i}\hat{I}_{t+h,i}-\hat{\gamma}_{t+h,i}\hat{I}_{t+h,i})|,\\ \mathcal{L}_{p}=\mathcal{L}_{d}(\hat{I}_{t+h,i})+\mathcal{L}_{o}(\hat{S}_{t+h,i},\hat{I}_{t+h,i},\hat{\beta}_{t+h,i},\hat{\gamma}_{t+h,i}),\end{gathered}$ (11)
where $x_{t+h,i}$ is the ground truth. $\mathcal{L}_{d}$ represents the error between predicted values and ground truth, which uses mean absolute error (MAE) as a loss measure. $\mathcal{L}_{o}$ is the residual term of the ODEs set. We use $\mathcal{L}_{d}$ to constrain the output of MLPs consistent with the real-world epidemic transmission dynamics and use $\mathcal{L}_{o}$ to constrain the output consistent with the mechanism of intrinsic epidemic transmission.
@@
where $\hat{y}_{t+h,i}\in\hat{\boldsymbol{Y}}_{t+h}$ is output value of HeatGNN and $x_{t+h,i}$ is the ground truth value. Finally, the main objective function $\mathcal{L}$ of HeatGNN is given as follows:
@@
where the loss term $\mathcal{L}_{g}$ is forecasting the error of our proposed framework and minimizing it to achieve good performance. Another loss term $\mathcal{L}_{p}$ is used as a regularizer of the EIEL module to enhance the ability of this module to represent the intrinsic epidemic transmission mechanism. The loss weight $\lambda$ is used to balance two different loss terms.
IV Experiments
In the experiment, we evaluate our proposed framework HeatGNN by answering the below research questions (RQs):
  • RQ1: Does HeatGNN outperform the existing state-of-the-art epidemic forecasting models?
  • RQ2: What is the impact of each core component in HeatGNN?
  • RQ3: What is the performance of HeatGNN with different hyperparameters?
  • RQ4: Is HeatGNN reliable for short-term and long-term epidemic forecasting?
  • RQ5: Does HeatGNN have robust scalability?
IV-A Datasets
We evaluate the performance of our framework using three real-world datasets following previous work[4]: Japan-Prefectures, US-States, and US-Regions.

### v2

@@ IV-B Epidemiology-informed Embedding Learning Module
The STGL module only captures the spatio-temporal heterogeneity among different locations, while many previous approaches predict epidemic transmission directly based on the learned ST embedding. However, only capturing the spatio-temporal pattern of the infected population is noisy for epidemic forecasting and lacks the intrinsic epidemic transmission mechanism, which is prone to producing generalization errors in predictions [wan2024epidemiology]. Therefore, we use the epidemiology mechanistic model to enhance the process of learning location transmission dynamics to capture latent patterns and reflect well-defined disease dissemination mechanisms. The time-varying SIR model can effectively capture the dynamic outbreak influenced by various complex factors and is adaptable across different epidemics [bucyibaruta2023discrete, zelenkov2023analysis]. Given a location $i$, we assign the time-varying SIR model as the epidemiology mechanistic model, which can be described as following ordinary differential equations (ODEs):
$\left\{\begin{aligned} \frac{dS_{i}}{dt}&=-\beta_{i}S_{i}I_{i},\\ \frac{dI_{i}}{dt}&=\beta_{i}S_{i}I_{i}-\gamma_{i}I_{i},\\ \frac{dR_{i}}{dt}&=\gamma_{i}I_{i},\end{aligned}\right$. (3)
where $S_{i}$, $I_{i}$, and $R_{i}$ represent the number of susceptible, infected, and recovered populations in location $i$, respectively. $\beta_{i}$ and $\gamma_{i}$ are the time-dependent parameters, which mean the infection rate of the susceptible population and the recovery rate of the infected population in the location $i$.
Although the time-varying SIR model reflects the intrinsic epidemic transmission mechanism, it requires high computational costs and domain knowledge to predict its large-scale parameters simultaneously. In addition, the variables of the time-varying SIR model are heavily entangled, and it is difficult to focus only on predicting changes in the number of infected populations ($I$) in epidemic forecasting. These challenges motivate us to propose the EIEL module that uses node embeddings to parameterize the five time-varying SIR variables. The EIEL module uses five different multi-layer perceptrons (MLPs) to predict susceptible population number $S$, infected population number $I$, recovered population number $R$, infection rate $\beta$, and recovery rate $\gamma$ in the time-varying SIR model at each location and time step. Each MLP consists of a fully connected (FC) layer with $L$ layers and an output layer. Additionally, to enhance the embedding expressiveness of the EIEL module, we learn epidemiology-informed location embeddings based on the ST embedding $\boldsymbol{P}_{t}=\{\boldsymbol{p}_{t,1},\boldsymbol{p}_{t,2},\ldots,\boldsymbol{p}_{t,N}\}\in\mathbb{R}^{N\times D_{1}}$. Taking $\hat{S}_{t+h,i}$ prediction as an example, given the $i$-th location ST embedding $\boldsymbol{p}_{t,i}\in\mathbb{R}^{D_{1}}$, the susceptible population embedding $\boldsymbol{u}_{i}^{(S)}\in\mathbb{R}^{D_{2}}$ of epidemiology-informed location embedding in $i$-th location in time step $t$ can be computed by the FC layer $g$ with $L$ layers:
$\boldsymbol{u}_{i}^{(S)}=g\left(\boldsymbol{p}_{t,i};\{\boldsymbol{W}^{l},\boldsymbol{b}^{l},\sigma^{l}\}_{l=1}^{L}\right),$ (4)
where $\boldsymbol{W}^{l}$ is the weight matrix of the $l$-th hidden layer, $\boldsymbol{b}^{l}$ is its bias vector, and $\sigma^{l}$ is the activation function. Then, the epidemic transmission prediction $\hat{S}_{t+h,i}$ is:
$\hat{S}_{t+h,i}=\sigma^{out}(\boldsymbol{W}^{out}\boldsymbol{u}_{i}^{(S)}+\boldsymbol{b}^{out}),$ (5)
where $\boldsymbol{W}^{out}$, $\boldsymbol{b}^{out}$, and $\sigma^{out}(\cdot)$ are the weight matrix, bias vector, and activation function in the output layer of the MLP, respectively. Therefore, the other four parameters ($\hat{I}_{t+h,i},\hat{R}_{t+h,i},\hat{\beta}_{t+h,i},\hat{\gamma}_{t+h,i}$) are predicted using four other MLPs with the same structure, where the parameters of these five MLPs are not shared. Finally, we obtain the epidemiology-informed location embedding $\boldsymbol{h}_{t,i}\in\mathcal{R}^{5\times D_{2}}$ for location $i$ in time step $t$, which contains five parameter embeddings of time-varying SIR. Moreover, variants of epidemiology mechanistic models such as SEIR or SIRD can be integrated into HeatGNN by modifying the MLPs in EIEL to align with model states and parameters. However, these extensions require additional data, which is often unavailable in mainstream datasets. In Section IV-E, we introduce the optimization to ensure that the epidemiology-informed location embedding is consistent with the intrinsic epidemic transmission mechanism.
IV-C Transmission Graph Module
Although the EIEL effectively captures the intrinsic epidemic transmission mechanisms of individual locations, it does not account for the interactive relationships between them. To address this, we propose a time-varying mechanistic affinity graph (MAG) that can efficiently characterize the mechanistic heterogeneity across locations. Unlike the geographical graph, the MAG provides a comprehensive representation of how transmission mechanisms vary over time and across different locations, reflecting the mechanistic dependencies that influence epidemic transmission. Moreover, to fully capture the complex heterogeneity by leveraging the mechanistic dependencies within the MAG, we propose the Heterogeneous Transmission Graph Network (HTGN), which encodes the transmission (dis)similarity between locations into their embeddings.
Mechanistic Affinity Graph. The MAG $\widetilde{\boldsymbol{M}}_{t}\in\mathbb{R}^{N\times N}$ is calculated by the cosine similarity function $\phi$ between the epidemiology-informed location embeddings of different locations. The mechanistic similarity between locations is defined as follows:
$m_{ij}=\phi(\boldsymbol{h}_{t,i},\boldsymbol{h}_{t,j}),$ (6)
where the mechanistic similarity $m_{ij}\in\boldsymbol{M}_{t}$ represents the mechanistic dependence between locations $i$ and $j$ based on epidemiology-informed location embeddings ($\boldsymbol{h}_{t,i}$ and $\boldsymbol{h}_{t,j}$) of two locations at time step $t$. $\boldsymbol{M}_{t}$ is a symmetric matrix. Then, we normalize the similarity between different locations:
$\widetilde{\boldsymbol{M}}_{t}=\boldsymbol{D}_{t}^{-\frac{1}{2}}\boldsymbol{M}_{t}\boldsymbol{D}_{t}^{-\frac{1}{2}},$ (7)
where $\boldsymbol{D}_{t}$ is the degree matrix of $\boldsymbol{M}_{t}$. To focus on key mechanistic dependence, reduce redundant information, and improve the computational efficiency of GNN, we design an indicator function to sparsify the MAG $\widetilde{\boldsymbol{M}}_{t}\in[0,1]$. This sparsification is controlled by a threshold $\delta$, applied to each normalized similarity $\widetilde{m}_{ij}\in\widetilde{\boldsymbol{M}}_{t}$:
@@ IV-C Transmission Graph Module
Heterogeneous Transmission Graph Network. The HTGN integrates MAG with the architecture of Spatio-Temporal Graph Convolutional Networks (STGCN) [yu2017spatio]. The HTGN consists of two Transmission-Temporal convolutional (TT-Conv) blocks. Each TT-Conv includes two temporal gated convolution layers and one transmission graph convolution. Given the historical infection number $\boldsymbol{X}_{t-w+1:t}$ and MAG $\widetilde{\boldsymbol{M}}_{t}\in\mathbb{R}^{N\times N}$ in time step $t$, we learn the transmission embedding $\boldsymbol{C}_{t}\in\mathbb{R}^{N\times D_{3}}$ of mechanistic heterogeneity by two TT-Conv blocks. For the input $\boldsymbol{v}^{l}$ of TT-Conv block $l$, the output $\boldsymbol{v}^{l+1}$ is computed by:
$\boldsymbol{v}^{l+1}=\Gamma_{1}^{l}*_{\tau}\operatorname{ReLU}\left(\Theta^{l}*_{\mathcal{G}}\left(\Gamma_{0}^{l}*_{\tau}\boldsymbol{v}^{l}\right)\right),$ (9)
where $\Gamma_{0}^{l}$ and $\Gamma_{1}^{l}$ are upper and lower temporal kernels in block $l$, respectively. $\Theta^{l}$ is the spectral kernel of the graph convolution in block $l$. $_{\tau}$ and $_{\mathcal{G}}$ are temporal and graph convolution operators, respectively. $\operatorname{ReLU}(\cdot)$ represents the activation function in block $l$. The input of the first TT-Conv is $\boldsymbol{v}^{(0)}=\boldsymbol{X}_{t-w+1:t}^{\top}\in\mathbb{R}^{N\times w}$. Then, we learn the transmission embedding $\boldsymbol{C}_{t}=\{\boldsymbol{c}_{t,1},\boldsymbol{c}_{t,2},\ldots,\boldsymbol{c}_{t,N}\}\in\mathbb{R}^{N\times D_{3}}$ from two TT-Conv blocks.
IV-D Epidemic Forecasting Decoder Module
@@
Finally, by simultaneously capturing the spatio-temporal heterogeneity and mechanistic heterogeneity across different locations, our framework applies an FC layer to predict the infection number, providing the epidemic forecasting result $\hat{\boldsymbol{Y}}_{t+h}$ in time step $t+h$:
$\hat{\boldsymbol{Y}}_{t+h}=\boldsymbol{W}[\boldsymbol{P}_{t};\boldsymbol{C}_{t}]+\boldsymbol{b},$ (10)
@@
where $\boldsymbol{W}$ and $\boldsymbol{b}$ are the weight matrix and bias vector, respectively, $\boldsymbol{P}_{t}$ is the ST embedding learned from the STGL module and $\boldsymbol{C}_{t}$ is the transmission embedding learned from the TG module.
IV-E Optimization
@@
To ensure the epidemiology-informed location embeddings are consistent with the mechanism of the intrinsic epidemic transmission, we use the idea from the physics-informed neural networks (PINN) [raissi2017physics] to propose a mechanistic loss to inform the learning process in the EIEL module. The physics loss incorporates the underlying physical laws of the intrinsic epidemic transmission mechanism into the EIEL module. Specifically, our objective is to solve the time-varying SIR model using five MLPs by minimizing the multi-objective mechanistic loss $\mathcal{L}_{p}$. Given the prediction of five variables ($\hat{S}_{t+h,i}$, $\hat{I}_{t+h,i}$, $\hat{R}_{t+h,i}$, $\hat{\beta}_{t+h,i}$, and $\hat{\gamma}_{t+h,i}$) of the time-varying SIR model, the loss $\mathcal{L}_{p}$ can be described as follows:
$\begin{gathered}\mathcal{L}_{d}(\hat{I}_{t+h,i})=\frac{1}{N}\sum_{i=1}^{N}|\hat{I}_{t+h,i}-x_{t+h,i}|,\\ \mathcal{L}_{o}(\hat{S}_{t+h,i},\hat{I}_{t+h,i},\hat{\beta}_{t+h,i},\hat{\gamma}_{t+h,i})\\ =\frac{1}{N}\sum_{i=1}^{N}|\frac{d\hat{I}_{t+h,i}}{dt}-(\hat{\beta}_{t+h,i}\hat{S}_{t+h,i}\hat{I}_{t+h,i}-\hat{\gamma}_{t+h,i}\hat{I}_{t+h,i})|,\\ \mathcal{L}_{p}=\mathcal{L}_{d}(\hat{I}_{t+h,i})+\mathcal{L}_{o}(\hat{S}_{t+h,i},\hat{I}_{t+h,i},\hat{\beta}_{t+h,i},\hat{\gamma}_{t+h,i}),\end{gathered}$ (11)
where $x_{t+h,i}$ is the ground truth. $\mathcal{L}_{d}$ represents the error between predicted values and ground truth, which uses mean absolute error (MAE) as a loss measure. $\mathcal{L}_{o}$ is the time derivatives residual between the neural network output and the time-varying SIR model. We use $\mathcal{L}_{d}$ to constrain the output of MLPs to be consistent with the real-world epidemic transmission. As the epidemiology-informed embeddings enable the quantification of mechanistic heterogeneity among locations, $\mathcal{L}_{o}$ puts a constraint on learning the infection trend defined by the epidemic model, i.e., the infection number’s derivative w.r.t. time $\frac{d\hat{I}_{t+h,i}}{dt}$. To achieve this, $\mathcal{L}_{o}$ approximates the second sub-equation in Equation 3 with estimated epidemic model parameters. Guided by $\mathcal{L}_{o}$, HeatGNN can more effectively learn the epidemic transmission mechanism at each location, thereby facilitating the quantification of mechanistic heterogeneity with the subsequent dynamic MAG. Moreover, the constraints are not only the $I$ but also $S$, $\beta$, $\gamma$, and $R$, which are indirectly constrained via the loss $\mathcal{L}_{o}$.
@@
where $\hat{y}_{t+h,i}\in\hat{\boldsymbol{Y}}_{t+h}$ is the output of HeatGNN and $x_{t+h,i}$ is the ground truth value. Finally, the main objective function $\mathcal{L}$ of HeatGNN is given as follows:
@@
where the term $\mathcal{L}_{g}$ is the forecasting error, and the term $\mathcal{L}_{p}$ is used as a regularizer of the EIEL module to enhance the ability of this module to represent the intrinsic epidemic transmission mechanism. The loss weight $\lambda$ is used to balance loss terms.
V Experiments
In the experiment, we evaluate HeatGNN by addressing the following research questions (RQs): RQ1: Does HeatGNN outperform the existing state-of-the-art epidemic forecasting models? RQ2: What is the impact of each core component in HeatGNN? RQ3: What is the performance of HeatGNN with different hyperparameters? RQ4: Can the mechanistic similarity reveal intrinsic epidemic transmission across locations? RQ5: Does HeatGNN exhibit robust scalability and efficient inference performance? RQ6: Does the HeatGNN robustness under noisy and incomplete data scenarios?
V-A Experimental Settings
Datasets and Metrics. We evaluate the performance of our HeatGNN using four real-world datasets with multiple wave outbreaks following previous works [deng2020cola, xie2022epignn]: Japan-Prefectures, US-States, US-Regions, and Australia-COVID: